OpenStudy (alekos):

Looking for the integral of the following.... sinx/(1+sin^2x)

OpenStudy (anonymous):

integral sinx*1/1+sin^2x take u = sinx

OpenStudy (alekos):

but du/dx = cosx dx

zepdrix (zepdrix):

Hmm that causes some trouble with the cosine that shows up D: yah

OpenStudy (alekos):

du/dx = cosx hence dx = du/cosx so we cant substitute

OpenStudy (alekos):

i'm missing something?

OpenStudy (anonymous):

ok ok ok ok :D i will always think of converting it to complex number :3 it looks more neat hehehe

OpenStudy (anonymous):

missing what ?

OpenStudy (anonymous):

converrrrtttt yeahhhhhhh

OpenStudy (anonymous):

hoy wanna work on complex ?

OpenStudy (anonymous):

u will only have one branch cut

OpenStudy (alekos):

Chad. If we differentiate your answer we get sinxcosx/(1+sin^2x)

OpenStudy (anonymous):

ohh a bit cool , continue wih trig identety u will got something like (k sin 2 x / m ( cos 2 x))

OpenStudy (anonymous):

the numerator and the denomominator divided by cos^2 (x). the integral bocomes : int ( sec(x)tan(x)/(sec^2 (x) + tan^2 (x)) dx = int sec(x)tan(x)/(2sec^2 (x) - 1) u = sec(x) du = sec(x)tan(x) dx = int du/(2u^2 - 1)

OpenStudy (anonymous):

lavoshneeey

OpenStudy (anonymous):

i think partial fraction is next step ....

OpenStudy (anonymous):

we would use it in first place instead

OpenStudy (anonymous):

but to me that only sound need to be solved with complex assumtion

OpenStudy (anonymous):

hmm but if u make it bounded u will only get an answer of real number, trust me -.- do you know how complex work ? @chad123 it looks hmm something wrong , u assumed first sin x = x then where is dx ?

OpenStudy (anonymous):

yeah i stuffed up what im trying to do is let sinx = x

OpenStudy (anonymous):

we will then have integral x/x^2+1 dx

OpenStudy (anonymous):

but then u have to make cos x dx = dx

OpenStudy (anonymous):

so it will be x cos x / 1+ x^2 dx

OpenStudy (anonymous):

u dint get it right ? its wrong

OpenStudy (anonymous):

1/(2u^2 - 1) = A/(sqrt(2)u+1) + B/(sqrt(2)u-1) i got A = -1/2 and B = 1/2

OpenStudy (alekos):

Your on the right track tanjung but i get 1/(2u^2-1) = 1/2(u+1/sqrt2)(u-1/sqrt2)

OpenStudy (alekos):

then i get 1/2(A/(u+1/sqrt2) + B/(u-1/sqrt2)

OpenStudy (alekos):

and i get A= -1/sqrt2 B= 1/sqrt2

OpenStudy (anonymous):